OpenStudy (anonymous):

Prove by induction: Summation where j=1 to n-1: j equals (n(n-1)/2)

OpenStudy (anonymous):

\[\sum_{j=1}^{n-1}j=\frac{ n(n-1) }{2 }\]

OpenStudy (anonymous):

Here is what I have, but it seems way to elementary

OpenStudy (anonymous):

Assume j=2. We gather that 2(2-1)/2=2 thus making the equation true. Now we will assume that for all k which will be equal to (n-1) now prove that it is also true for k+1.

OpenStudy (anonymous):

\[\frac{ (k+2)(k+1) }{ 2 }=\frac{ k^2+3k+2 }{ 2 }\]

OpenStudy (anonymous):

If it works for n then you have for n+1: \[\sum_{k=1}^{n+1}k=\sum_{k=1}^{n-1}k + (n+1)+n = \frac{n(n-1)}{2}+2n+1 =\frac{n^2-n+4n+2}{2}=\frac{(n+1)(n+2)}{2}\] As you can rewrite: \[\sum_1^{n-1}k =\frac{n(n-1)}{2} \implies \sum_{k=1}^{n}\frac{n(n+1)}{2}\] So this transforms into: \[\frac{n+1((n+1)+1)}{2}\] Which is what you want.

OpenStudy (anonymous):

so in essence I was right I just had to separate the k+2, got it! thanks